# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ 2.2 Generalized Nonparametric Regression
$\widehat{f}^{*}_{n}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}\Big{[}-y_{i}f(\mathbf{x}_{i})+n_{\text{trial}}\log\{1+e^{f(\mathbf{x}_{i})}\}\Big{]}.$
2.3 Approximate Empirical Risk Minimization and Function Smoothness
As the DNN estimator (2.4) is an approximate empirical loss minimizer, we use the following $\Delta_{n}(\widehat{f}_{n},f_{0})$ to measure the difference between the expected empirical risk of $\widehat{f}_{n}$ and the global minimum over all networks in the class, that is,
@@ 2.3 Approximate Empirical Risk Minimization and Function Smoothness
$\Delta_{n}(\widehat{f}_{n},f_{0})=\mathbb{E}_{f_{0}}\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}-y_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\psi(\widehat{f}_{n}(\mathbf{x}_{i}))-\inf_{f\in\mathcal{F}(L,\mathbf{p},s,\mathcal{F})}\frac{1}{n}\sum_{i=1}^{n}-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i}))\bigg{\}}.$ (2.5)
Obviously, $\Delta_{n}(\widehat{f}_{n},f_{0})=0$ if $\widehat{f}_{n}$ is an empirical risk minimizer. We also define
$\ell(\mathbf{x};\widehat{f}_{n},f_{0})=-\psi^{\prime}(f_{0}(\mathbf{x}))\widehat{f}_{n}(\mathbf{x})+\psi(\widehat{f}_{n}(\mathbf{x}))+\psi^{\prime}(f_{0}(\mathbf{x}))f_{0}(\mathbf{x})-\psi(f_{0}(\mathbf{x})),$
which measures the discrepancy between the predictor $\widehat{f}_{n}$ and the true function $f_{0}$ at a point $\mathbf{x}$ outside the training sample, $(\mathbf{x}_{i},y_{i})_{i=1}^{n}$. The prediction error $R_{n}(\widehat{f}_{n},f_{0})$ is then defined as the expected value of the loss function:
$R_{n}(\widehat{f}_{n},f_{0}):=\mathbb{E}[\ell(\bm{X};\widehat{f}_{n},f_{0})],$ (2.6)
where $\bm{X}$ is an out of sample point such that $\bm{X}\stackrel{{\scriptstyle\mathcal{D}}}{{=}}\mathbf{x}_{1}$ and is independent of $(\mathbf{x}_{i},y_{i})_{i=1}^{n}$.
For desirable properties of $\widehat{f}_{n}$, we assume that $f_{0}$ belongs to a Hölder class of smooth functions. This assumption provides a flexible framework for characterizing function smoothness, which is important for establishing convergence rates, consistency, and estimator optimality. It allows the use of approximation theory and functional analysis to derive meaningful statistical guarantees. Widely adopted in the literature (Schmidt-Hieber, 2020; Zhong et al., 2022; Fan and Gu, 2024), the Hölder class of smooth functions with parameters $\gamma$, $K>0$, and domain $\mathbb{D}\subset\mathbb{R}^{r}$ is
@@
where $\lfloor\gamma\rfloor$ is the largest integer strictly smaller than $\gamma$, and $\partial^{{\bm{\beta}}}=\partial^{\beta_{1}}\ldots\partial^{\beta_{r}}$ with ${\bm{\beta}}=(\beta_{1},...,\beta_{r})^{\top}\in\mathbb{N}^{r}$. We further assume that the nonparametric function $f_{0}$ is a $(q+1)$-composition ($q\in\mathbb{N}$) of several functions, that is for some vectors $\mathbf{d}=(d_{0},...,d_{q+1})\in\mathbb{N}_{+}^{q+2}$, $\mathbf{t}=(t_{0},...,t_{q})\in\mathbb{N}_{+}^{q+1}$ and ${\bm{\gamma}}=(\gamma_{0},...,\gamma_{q})\in\mathbb{R}_{+}^{q+1}$, $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$, where
[…]
$g_{ij}\in\mathcal{G}_{t_{i}}^{\gamma_{i}}([a_{i},b_{i}]^{t_{i}})\text{ with }|a_{i}|,|b_{i}|\leq K\bigg{\}},$
and $\mathbf{d}$ and $\mathbf{t}$ characterize the dimensions of the function, while ${\bm{\gamma}}$ represents the measure of smoothness. For instance, if
@@
and $g_{ij}$ are twice continuously differentiable, then smoothness ${\bm{\gamma}}=(2,2,2)$, dimensions $\mathbf{d}=(24,6,3,1)^{\top}$ and $\mathbf{t}=(4,2,3)^{\top}$. We further define $\gamma_{j}^{*}=\gamma_{j}\Pi_{l=i+1}^{q}(\gamma_{l}\wedge 1)$ and $\phi_{n}=\max_{j=0,...,q}n^{-\frac{2\gamma_{j}^{*}}{2\gamma_{j}^{*}+t_{j}}}$.
@@ 2.4 Ensemble Subsampling Method (ESM)
Consider a scenario where $\mathbf{x}$ is drawn from an unknown probability measure $\mathbb{P}_{\mathbf{x}}$ supported on $\mathcal{X}$, and for an input $\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*}\in\mathcal{X}$, the unobserved outcome $y_{{\mathop{\mathrm{new}}}}$ follows (2.3). Our goal is to predict the conditional expectation $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*})=\psi^{\prime}(f_{0}(\mathbf{x}_{*}))$ along with its uncertainty. For this, we introduce the Ensemble Subsampling Method (ESM), based on subsampling techniques (Wager and Athey, 2018; Fei and Li, 2024), to construct ensemble predictors and confidence intervals. ESM consists of the following components:
  • (Subsampling) Let =$\mathcal{I}=\{1,\dots,n\}$ denote the index set of the training dataset $\mathcal{D}_{n}$. We construct subsets of size $r(<n)$, yielding $B^{*}=\binom{n}{r}$ unique combinations. Denoting the $b$-th subset as $\mathcal{I}^{b}=\{i_{1},\dots,i_{r}\}$, where $i_{1}<\dots<i_{r}$ and $b\in[1:B^{*}]$.
  • (Predictor Construction) For each subset $\mathcal{I}^{b}$, the model is trained using the observations indexed by $\mathcal{I}^{b}$. The corresponding predictor $\widehat{f}^{b}\in\mathcal{F}(L,\mathbf{p},s,F)$ is obtained by solving the optimization problem:
    $\widehat{f}^{b}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{|\mathcal{I}^{b}|}\sum_{i\in\mathcal{I}^{b}}-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i})).$
  • (Ensemble Prediction) Using the predictors $\{\widehat{f}^{b}:b=1,\dots,B^{*}\}$, we compute the ensemble prediction for any $\mathbf{x}_{*}\in\mathcal{X}$ as:
@@
    The conditional expectation is then estimated as $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*})=\psi^{\prime}(\widehat{f}^{B^{*}}(\mathbf{x}_{*})))$.
@@
    $\text{CI}(\mathbf{x}_{*})=\big{[}\psi^{\prime}(\widehat{f}^{B^{*}}(\mathbf{x}_{*})-c_{\alpha}\widehat{\sigma}_{*}),\psi^{\prime}(\widehat{f}^{B^{*}}(\mathbf{x}_{*})+c_{\alpha}\widehat{\sigma}_{*})\big{]}.$ (2.7)
@@
The distribution of $\widehat{f}^{b}(\mathbf{x}_{*})$ for a fixed $\mathbf{x}_{*}$ is intractable. However, since $\widehat{f}^{b}(\mathbf{x}_{*})$ is permutation symmetric with respect to $\mathcal{I}^{b}$, the ensemble predictor $\widehat{f}^{B^{*}}(\mathbf{x}_{*})$ forms a generalized U-statistic, which we later show to be asymptotically normal by leveraging U-statistics theory. In practice, we allow $r$ to grow with $n$, but evaluating all $B^{*}={{n}\choose{r}}$ models becomes computationally infeasible. To overcome this, we use a stochastic approximation by randomly drawing $B$ size-$r$ subsamples from $\mathcal{D}_{n}$. This involves independently sampling indices $b_{1},\dots,b_{B}$ from $1$ to (${{n}\choose{r}}$, where each subsample is indexed by $\mathcal{I}^{b_{j}}$. The estimator $\widehat{f}^{B}(\mathbf{x}_{*})$ is defined as
@@
As (2.8) provides the predictor for $f_{0}$, the variance $\widehat{\sigma}_{*}$ in (2.7) in the Confidence Interval Construction procedure is given as
$\widehat{\sigma}_{*}^{2}=\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\widehat{V}_{i}^{2}.$ (2.9)
Here, $\widehat{V}_{i}$ is defined as
$\widehat{V}_{i}=\frac{\sum_{j=1}^{B}(J_{b_{j}i}-J_{\cdot i})(\widehat{f}^{b_{j}}(\mathbf{x}_{*})-\widehat{f}^{B}(\mathbf{x}_{*}))}{B},$
where $J_{b_{j}i}=I(i\in\mathcal{I}^{b_{j}})$ and $J_{\cdot i}=\frac{1}{B}\sum_{j=1}^{B}J_{b_{j}i}$.
@@ 3 Theoretical Results
We study the asymptotic properties of the ensemble predictor $\widehat{f}_{n}^{B}$ in (2.8). We first impose some regularity assumptions.
@@
Suppose that (2.3) holds. The support of $y$, denoted by $\mathrm{supp}(h(y))$ is fixed and does not depend on $f_{0}(\mathbf{x})$. In other words, $h(y)>0$ for all $y\in\mathrm{supp}(h)$. Moreover, the sufficient statistic $T(y)=y$ is not degenerate and can take at least two distinct values in $\mathrm{supp}(h)$.
@@
For any $\mathbf{x}$, the conditional distribution $p(y|\bm{X}=\mathbf{x})$ is sub-exponential. Specifically, there exists a universal constant $\kappa>0$ such that for all $t>0$,
$\mathbb{P}(|y-\mathbb{E}[y|\bm{X}=\mathbf{x}]|\geq t|\bm{X}=\mathbf{x})\leq 2\exp\left(-\frac{t}{\kappa}\right).$
@@
The network class $\mathcal{F}(L,\mathbf{p},s,F)$ with $L$, $\mathbf{p}$ and $s$ satisfies
  1. 1.
    $F\geq\max(K,1)$,
  2. 2.
    $\sum_{i=0}^{q}\log_{2}(4t_{i}\vee 4\gamma_{i})\log_{2}n\leq L\precsim\log^{\alpha}n$ for some $\alpha>1$,
  3. 3.
    $n\phi_{n}\precsim\min_{i=1,...,L}p_{i}\leq\max_{i=1,...,L}p_{i}\precsim n$,
  4. 4.
    $s\precsim n\phi_{n}\log n$.
Assumption 3.1 is mild, ensuring that $p(y|\bm{X}=\mathbf{x})$ remains well-defined and avoids collapsing into a degenerate distribution. Additionally, that $T(y)=y$ takes at least two distinct values in $\mathrm{supp}(h)$ guarantees a non-zero variance of $y$, ensuring the log-partition function $\psi(f_{0}(\mathbf{x}))$ exhibits local convexity. Assumption 3.2 is technical and ensures the concentration of the label variable $Y$ does not deviate too much. Assumptions 3.1 and 3.2 are satisfied by most common exponential family distributions, such as $\chi^{2}_{K}$ with fixed $K$, Beta, Bernoulli, Poisson, and Gaussian distributions, making them broadly applicable. Assumption 3.3 is a technical assumption, which specifies the requirements on the network parameters relative to the sample size $n$ and smoothness parameters $\phi_{n}$, $t_{i}$ and $\gamma_{i}$, including the number of layers $L$ and the sparsity $s$.
We let $\mathbf{z}_{i}=(y_{i},\mathbf{x}_{i})$ represent an independent observation of sample points. To further investigate the asymptotic properties of $\widehat{f}_{n}^{B}$, we introduce
$\xi_{1,r}(\mathbf{x})=\operatorname{\rm Cov}(\widehat{f}(\mathbf{x};\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{r}),\widehat{f}(\mathbf{x},\mathbf{z}_{1},\mathbf{z}^{\prime}_{2},...,\mathbf{z}^{\prime}_{r})),$
where $\mathbf{x}\in\mathcal{X}$ is in the support of probability measure $\mathbb{P}_{\mathbf{x}}$, and $\xi_{1,r}$ quantifies the covariance between predictions based on two overlapping subsets of size $r$ that share a common point $\mathbf{z}_{1}$ but differ in the remaining elements. This term represents a second-order component in the Hoeffding decomposition of a U-statistic (Hoeffding, 1992), capturing the pairwise dependence structure within the subsamples used for constructing the estimator. Here, $\widehat{f}(\mathbf{x};\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{r})$ means that we obtain $\widehat{f}$ from the subsample $\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{r}$, and then apply it to the point $\mathbf{x}$. The terms $\mathbf{z}^{\prime}_{i}$ and $\mathbf{z}_{i}$ are independently generated from the same data generation process.
@@
The covariance $\xi_{1,r}(\mathbf{x})$ for any given $\mathbf{x}$ satisfies $\lim_{n\to\infty}1/(n\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x}))=0$. The sub sampling size $r$ satisfies $\lim_{n\to\infty}r/n=0$ and $\lim_{n\to\infty}n\phi_{r}L\log^{2}(r)/(r^{2}\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x}))=0$. Here, $\phi_{r}=\max_{j=0,...,q}r^{-(2\gamma_{j}^{*})/(2\gamma_{j}^{*}+t_{j})}$. Furthermore, the number of subsampling times $B$ is large such that $\lim_{n\to\infty}n/B=0$.
Assumption 3.4 gives assumptions on the covariance $\xi_{1,r}(\mathbf{x})$, the subsample size $r$, total sample size $n$, and number of subsampling repetitions $B$. It allows some degeneracy in the first-order U-statistic covariance but ensures its rate remains above $n^{-1}$ to keep the first-order contribution identifiable. The condition $\lim_{n\to\infty}n\phi_{r}L\log^{2}(r)/(r^{2}\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x}))=0$ is a necessary lower bound on $r$ to guarantee sufficient training information per subsample. It indicates that a sufficiently large $r$ is required to ensure deep neural networks learn $f_{0}$ effectively; one example is $r=n\phi_{n}$. A large enough $B$ minimizes subsampling noise, preserving the estimator’s asymptotic behavior and accuracy by keeping variability from limited subsamples negligible.

### v3

@@ 2.2 Generalized Nonparametric Regression
$\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}=\operatorname{argmin}_{f\in\mathcal{F}(L,\mathbf{p},s,F)}\frac{1}{n}\sum_{i=1}^{n}\Big[-y_{i}f(\mathbf{x}_{i})+n_{\text{trial}}\log\{1+e^{f(\mathbf{x}_{i})}\}\Big].$
2.3 Empirical Risk Minimization and Function Smoothness
Recall that the pointwise Kullback–Leibler divergence between two distributions of the form (2.3), one specified by $f$ and another by the true $f_{0}$, conditional on $\mathbf{x}$, i.e.,
$\mathbb{E}\left[-yf(\mathbf{x})+\psi(f(\mathbf{x}))+yf_{0}(\mathbf{x})-\psi(f_{0}(\mathbf{x}))\,\big|\,\mathbf{x}\right]=-\psi^{\prime}(f_{0}(\mathbf{x}))(f(\mathbf{x})-f_{0}(\mathbf{x}))+\psi(f(\mathbf{x}))-\psi(f_{0}(\mathbf{x})),$
where we use $\mathbb{E}[y|\mathbf{x}]=\psi^{\prime}(f_{0}(\mathbf{x}))$. This equals the Bregman divergence between $f(\mathbf{x})$ and $f_{0}(\mathbf{x})$, which is nonnegative and vanishes if and only if $f(\mathbf{x})=f_{0}(\mathbf{x})$. We then quantify the estimation error of $\widehat{f}$, any estimator of $f_{0}$, by evaluating the pointwise estimation error at an independent test point $\bm{X}\sim\mathbf{x}_{1}$: $\ell(\bm{X};\widehat{f},f_{0})=-\psi^{\prime}(f_{0}(\bm{X}))(\widehat{f}(\bm{X})-f_{0}(\bm{X}))+\psi(\widehat{f}(\bm{X}))-\psi(f_{0}(\bm{X})).$ We further define the average estimation error, our main theoretical target, as
$R_{n}(\widehat{f},f_{0}):=\mathbb{E}[\ell(\bm{X};\widehat{f},f_{0})],$ (2.5)
where the expectation is taken with respect to all involved random variables.
Computing the exact minimizer $\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}$ in (2.4) is challenging due to the nonconvex loss (Fan & Gu, 2024). Instead, we obtain an approximate minimizer $\widehat{f}_{n}$ using optimization algorithms such as gradient descent. To quantify the optimization error, we introduce
@@ 2.3 Empirical Risk Minimization and Function Smoothness
$\Delta_{n}(\widehat{f}_{n},\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}})=\mathbb{E}\bigg\{\frac{1}{n}\sum_{i=1}^{n}(-y_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\psi(\widehat{f}_{n}(\mathbf{x}_{i}))-\frac{1}{n}\sum_{i=1}^{n}(-y_{i}\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}(\mathbf{x}_{i})+\psi(\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}(\mathbf{x}_{i}))\bigg\}$ (2.6)
that measures the difference between the expected empirical risk of $\widehat{f}_{n}$ and $\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}$. For notational ease, we write $\Delta_{n}(\widehat{f}_{n}):=\Delta_{n}(\widehat{f}_{n},\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}})$ when the reference to $\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}$ is clear. We will show that a small empirical deviation $\Delta_{n}(\widehat{f}_{n})$, combined with controlled model complexity of the neural network class, implies a bound on the average excess risk $R_{n}(\widehat{f}_{n},f_{0})$, forming the foundation for the inference theory developed in subsequent sections.
For desirable properties of $\widehat{f}_{n}$, we assume $f_{0}$ belongs to a Hölder class (Schmidt-Hieber, 2020; Fan & Gu, 2024) with parameters $\gamma$, $K>0$, and domain $\mathbb{D}\subset\mathbb{R}^{r}$:
@@
where $\lfloor\gamma\rfloor$ is the largest integer strictly smaller than $\gamma$, and $\partial^{{\bm{\beta}}}=\partial^{\beta_{1}}\ldots\partial^{\beta_{r}}$ with ${\bm{\beta}}=(\beta_{1},...,\beta_{r})^{\top}\in\mathbb{N}^{r}$. We further assume that $f_{0}$ is a $(q+1)$-composition ($q\in\mathbb{N}$) of several Hölder functions. That is, for some vectors $\mathbf{d}=(d_{0},...,d_{q+1})\in\mathbb{N}_{+}^{q+2}$, $\mathbf{t}=(t_{0},...,t_{q})\in\mathbb{N}_{+}^{q+1}$ and ${\bm{\gamma}}=(\gamma_{0},...,\gamma_{q})\in\mathbb{R}_{+}^{q+1}$, $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$, where
$\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K):=\bigg\{$ $f=g_{q}\circ\cdots\circ g_{0}:g_{i}=(g_{ij})_{j}:[a_{i},b_{i}]^{d_{i}}\to[a_{i+1},b_{i+1}]^{d_{i+1}},$
$g_{ij}\in\mathcal{G}_{t_{i}}^{\gamma_{i}}([a_{i},b_{i}]^{t_{i}})\text{ with }|a_{i}|,|b_{i}|\leq K\bigg\}.$
The motivation for the assumptions on $f_{0}$ is that the Hölder class provides a flexible framework for quantifying smoothness, capturing both integer and fractional orders, and serves as a natural benchmark for assessing neural network approximation (Yarotsky, 2017), while the compositional structure mirrors the architectural design of deep neural networks with each layer applying a transformation to the output of the previous one (Yarotsky, 2017). We also note that $\mathbf{d}$ and $\mathbf{t}$ characterize the dimensions of the function, while $\bm{\gamma}$} represents the measure of smoothness. For instance, if
@@
and $g_{ij}$ are twice continuously differentiable, then smoothness ${\bm{\gamma}}=(2,2,2)$, dimensions $\mathbf{d}=(24,6,3,1)^{\top}$ and $\mathbf{t}=(4,2,3)^{\top}$. We further define $\gamma_{j}^{*}=\gamma_{j}\Pi_{l=i+1}^{q}(\gamma_{l}\wedge 1)$ and $\phi_{n}=\max_{j=0,...,q}n^{-2\gamma_{j}^{*}/(2\gamma_{j}^{*}+t_{j})}$.
@@ 2.4 Ensemble Subsampling Method (ESM)
Consider a scenario where $\mathbf{x}$ is drawn from an unknown probability measure $\mathbb{P}_{\mathbf{x}}$ supported on $\mathcal{X}$, and for an input $\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*}\in\mathcal{X}$, the unobserved outcome $y_{{\mathop{\mathrm{new}}}}$ follows (2.3). Our goal is to estimate $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*})=\psi^{\prime}(f_{0}(\mathbf{x}_{*}))$ and its uncertainty. For this, we introduce the Ensemble Subsampling Method (ESM), based on subsampling techniques (Wager & Athey, 2018; Fei & Li, 2024), to construct ensemble estimators and confidence intervals. ESM consists of the following components, which are summarized in Figure 1 as well.
  • (Subsampling) Let $\mathcal{I}=\{1,\dots,n\}$ denote the index set of the training dataset $\mathcal{D}_{n}$. We construct subsets of size $r(<n)$, yielding $B^{*}=\binom{n}{r}$ unique combinations. Denoting the $b$-th subset as $\mathcal{I}^{b}=\{i_{1},\dots,i_{r}\}$, where $i_{1}<\dots<i_{r}$ and $b\in[1:B^{*}]$.
  • (Estimator Construction) For each $1\leq b\leq B^{*}$, we aim to minimize the objective function in (2.4) within $\mathcal{F}(L,\mathbf{p},s,F)$ by using a gradient based optimization algorithm such as Stochastic Gradient Descent (SGD) on the observations indexed by $\mathcal{I}^{b}$. Let $\widehat{f}^{b}$ denote the resulting approximate minimizer satisfying
    $\Delta_{r}(\widehat{f}^{b},\widehat{f}^{b,{\mathop{\mathrm{opt}}}})\leq\Delta^{{\mathop{\mathrm{opt}}}}_{b},$ (2.7)
    where $\Delta^{{\mathop{\mathrm{opt}}}}_{b}$ quantifies the optimization error (Fan & Gu, 2024), and $\widehat{f}^{b,{\mathop{\mathrm{opt}}}}$ and $\Delta_{r}(\widehat{f}^{b},\widehat{f}^{b,{\mathop{\mathrm{opt}}}})$ are defined respectively in (2.4) and (2.6) on the subsample indexed by $\mathcal{I}^{b}$.
  • (Ensemble Estimation) Using the estimators $\{\widehat{f}^{b}:b=1,\dots,B^{*}\}$, we compute the ensemble estimation for any $\mathbf{x}_{*}\in\mathcal{X}$ as:
@@
    and estimate $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*})$ as $\psi^{\prime}(\widehat{f}^{B^{*}}(\mathbf{x}_{*})))$.
@@
[…]
@@
Figure 1: Overview of the Ensemble Subsampling Method (ESM).
The distribution of $\widehat{f}^{b}(\mathbf{x}_{*})$ for a fixed $\mathbf{x}_{*}$ is intractable. However, since $\widehat{f}^{b}(\mathbf{x}_{*})$ is permutation symmetric with respect to $\mathcal{I}^{b}$, the ensemble estimator $\widehat{f}^{B^{*}}(\mathbf{x}_{*})$ forms a generalized U-statistic, which we later show to be asymptotically normal by leveraging U-statistics theory. In practice, we allow $r$ to grow with $n$, but evaluating all $B^{*}={\binom{n}{r}}$ models becomes computationally infeasible. To overcome this, we use a stochastic approximation by randomly drawing $B$ size-$r$ subsamples from $\mathcal{D}_{n}$. This involves independently sampling indices $b_{1},\dots,b_{B}$ from $1$ to $\binom{n}{r}$}, where each subsample is indexed by $\mathcal{I}^{b_{j}}$, and computing
@@
As (2.9) provides the estimator for $f_{0}$, the variance $\widehat{\sigma}_{*}$ in (2.8) in the Confidence Interval Construction procedure is given as
$\widehat{\sigma}_{*}^{2}=\underbrace{\frac{n(n-1)}{(n-r)^{2}}\sum_{i=1}^{n}\widehat{V}_{i}^{2}}_{\text{original }\widehat{\sigma}^{2}}-{\underbrace{\frac{n(n-1)}{(n-r)^{2}}\cdot\frac{1}{B(B-1)}\sum_{i=1}^{n}\sum_{j=1}^{B}(Z_{b_{j}i}-\widehat{V}_{i})^{2}}_{\text{bias correction}}}.$ (2.10)
Here, $Z_{b_{j}i}$ and $\widehat{V}_{i}$ are defined as
$Z_{b_{j}i}=(J_{b_{j}i}-J_{\cdot i})\big(\widehat{f}^{b_{j}}(\mathbf{x}_{*})-\widehat{f}^{B}(\mathbf{x}_{*})\big),\quad\widehat{V}_{i}=\frac{\sum_{j=1}^{B}Z_{b_{j}i}}{B},$
where $J_{b_{j}i}=I(i\in\mathcal{I}^{b_{j}})$ and $J_{\cdot i}=\frac{1}{B}\sum_{j=1}^{B}J_{b_{j}i}$. Extending Peng et al. (2024), (2.10) jointly corrects Monte Carlo bias and finite-sample effects using the factor $n(n-1)/(n-r)^{2}$, enabling valid inference with $B\succsim n$. Its construction relies on the dominance of the first order term in the Hoeffding decomposition, and the bias correction term adequately captures its resampling variability, regardless of the specific scaling between $r$ and $n$.
@@ 3 Theoretical Results
Let $\mathbf{z}_{i}=(y_{i},\mathbf{x}_{i})$ represent an independent observation of sample points. For studying the asymptotic properties of $\widehat{f}^{B}$ in (2.9), we introduce
$\xi_{1,r}(\mathbf{x})=\operatorname{\rm Cov}(\widehat{f}(\mathbf{x};\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{r}),\widehat{f}(\mathbf{x};\mathbf{z}_{1},\mathbf{z}^{\prime}_{2},...,\mathbf{z}^{\prime}_{r})).$ (3.1)
Here $\mathbf{x}\in\mathcal{X}$ is in the support of probability measure $\mathbb{P}_{\mathbf{x}}$, $\widehat{f}(\mathbf{x};\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{r})$ means that we obtain $\widehat{f}$ from the subsample $\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{r}$ and then apply it to the point $\mathbf{x}$, and $\xi_{1,r}$ quantifies the covariance between estimates based on two overlapping subsets of size $r$ that share a common point $\mathbf{z}_{1}$ but differ in the remaining elements. $\xi_{1,r}$ represents a second-order component in the Hoeffding decomposition of a U-statistic (Hoeffding, 1992), capturing the pairwise dependence structure within the subsamples used for constructing the estimator. The terms $\mathbf{z}^{\prime}_{i}$ and $\mathbf{z}_{i}$ are independently generated from the same data generation process. We impose some regularity assumptions for our theoretical guarantees.
@@
Suppose that (2.3) holds and the true function $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$. The support of $y$, denoted by $\mathrm{supp}(h(y))$, is fixed and does not depend on $\mathbf{x}$. Moreover, the sufficient statistic $T(y)=y$ is not degenerate and can take at least two distinct values in $\mathrm{supp}(h)$. For any $\mathbf{x}$, the conditional distribution $p(y|\mathbf{x})$ is sub-exponential. Specifically, with bounded $f_{0}$ there exists a universal constant $\kappa>0$ such that for all $t>0$,
$\mathbb{P}(|y-\mathbb{E}(y|\mathbf{x})|\geq t|\mathbf{x})\leq 2\exp\left(-\frac{t}{\kappa}\right).$
@@
The network class $\mathcal{F}(L,\mathbf{p},s,F)$ with $L$, $\mathbf{p}$ and $s$ satisfies i) $F\geq\max(K,1)$, ii) $\sum_{i=0}^{q}\log_{2}(4t_{i}\vee 4\gamma_{i})\log_{2}n\leq L\precsim\log^{\alpha}n$ for some $\alpha>1$, iii) $n\phi_{n}\precsim\min_{i=1,...,L}p_{i}\leq\max_{i=1,...,L}p_{i}\precsim n$, iv) $s\precsim n\phi_{n}\log n$. Here, $\phi_{n}=\max_{j=0,...,q}n^{-2\gamma_{j}^{*}/(2\gamma_{j}^{*}+t_{j})}$, where $\gamma_{j}^{*}=\gamma_{j}\Pi_{l=i+1}^{q}(\gamma_{l}\wedge 1)$. The values $t_{i},\gamma_{i}$ and $p_{i}$ are the $(i+1)$-th element in the vectors $\mathbf{t}$, $\bm{\gamma}$} and $\mathbf{p}$, respectively.
@@
For each subsample with index $\mathcal{I}^{b}$ and size $r$, there exists a constant $C>0$ such that the optimization error $\Delta^{{\mathop{\mathrm{opt}}}}_{b}$ defined in (2.7) satisfies $\Delta^{{\mathop{\mathrm{opt}}}}_{b}\leq C\phi_{r}L\log^{2}r$.
@@
The covariance $\xi_{1,r}(\mathbf{x})$ defined in (3.1) satisfies $\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x})\geq Cr^{-(1+\varepsilon)}$ for some small $\varepsilon>0$. The subsample size $r=n^{\gamma}$ satisfies $1/(1+\min_{j=0,...,q}(2\gamma_{j}^{*})/(2\gamma_{j}^{*}+t_{j}))<\gamma<1$, and the number of subsampling times $B$ is large such that $B\succsim n$.
Assumption 3.1 ensures that $p(y|\mathbf{x})$ not degenerate, while the concentration of $y$ does not deviate too widely. Additionally, that $T(y)=y$ takes at least two distinct values in $\mathrm{supp}(h)$ guarantees a non-zero variance of $y$, ensuring the log-partition function $\psi(f_{0}(\mathbf{x}))$ exhibits local convexity. The boundedness of $f_{0}$ guarantees the existence of $\kappa$. This condition, weaker than the sub-Gaussian tail bound assumed in Fan et al. (2024), is satisfied by most common exponential family distributions, including Beta, Bernoulli, Poisson, and Gaussian. Assumption 3.2 is a technical assumption similar to Schmidt-Hieber (2020), which specifies the requirements on the network parameters relative to the sample size $n$ and smoothness parameters $\phi_{n}$, $t_{i}$ and $\gamma_{i}$, including the number of layers $L$ and the sparsity $s$. Assumption 3.3, related to the training dynamics of DNNs, stipulates that DNNs return an estimator sufficiently close to the global minimum. We note that the rate in Assumption 3.3 pertains to the prediction error in Theorem 3.5. Our intent is not to assert an explicit convergence rate for SGD in practice, but to impose a sufficient condition that separates optimization error from the inferential analysis. This perspective follows standard practice in the statistical literature (Chernozhukov et al., 2018; Schmidt-Hieber, 2020; Fan & Gu, 2024), where such assumptions serve as technical devices to ensure valid inference rather than as primary objects of study. Assumption 3.4 on $\xi_{1,r}(\mathbf{x})$ is not meant to serve as a precise characterization of neural networks. Rather, it ensures that the first-order Hoeffding projection remains non-degenerate and dominates the higher-order components, as in the random forest setting (Wager & Athey, 2018). We choose the subsample size $r$ much smaller than $n$ to ensure first-order dominance but large enough to limit bias, and let the number of subsamples $B$ scale with $n$ to control computation.